The lengths (in inches) of two sides of a regular pentagon are represented by the expressions 5x - 27 and 2x - 6 . find the length of a side of the pentagon
Answer by mananth(16949)

5x-27 = 2x-6
why?
because the sides of regular pentagon are equal
5x-2x=-6+27
3x=21
x= 21/3
x=7
..
2*7-6
14-6 =8 inches